OpenStudy (anonymous):

Solve the system of equations. 6x + 2y = 10 and -6x - 2y = 12 A. (0, 5) B. (-1, 8) C. (2, -1) D. no solution E. infinitely many solutions

OpenStudy (theviper):

D.no solution

OpenStudy (anonymous):

TheViper is correct because it comes out to 0=22

OpenStudy (anonymous):

Since both the xs and ys cancel out

OpenStudy (theviper):

\[\frac{6}{-6}=\frac{2}{-2} \neq \frac{-10}{-12}\]

OpenStudy (anonymous):

not possible